# Measurement incompatibility in Bayesian multiparameter quantum estimation

### v1

@@ Acknowledgements.
FA acknowledges fruitful discussions with Stefano Chessa, Rafał Demkowicz-Dobrzański, Vittorio Giovannetti and Ludovico Lami. FA acknowledges partial financial support from Marie Skłodowska-Curie Action EUHORIZON-MSCA-2021PF-01 (project QECANM, Grant No. 101068347). DB acknowledges support from the European Union’s Horizon Europe research and innovation program under the EPIQUE Project (Grant Agreement No. 101135288). JR acknowledges financial support from the Surrey Future Fellowship Programme.
@@ Appendix A Prior loss and precision gain in Bayesian quantum estimation
where $\lambda$ was already defined in Eq. (15). Such an optimal prior loss is is nothing but $\operatorname{tr}[L\Sigma_{p}]$, where $\Sigma_{p}$ is the covariance matrix of the multivariate prior probability $p(\boldsymbol{\theta})$ and we have introduced a lowercase trace of $d{\times}d$ matrices, distinct from the uppercase trace of operators acting on the physical Hilbert space of the system.
@@
Recent literature has referred to the quantity $L^{ij}\operatorname{Tr}[\rho_{0}S_{j}S_{i}]$ as an alternative definition of precision gain when the SPM bound is used as an optimiser for control parameters [overton2024five, rubio2025on]. From a computational standpoint, this definition is often more convenient, and it has likewise been adopted in our code [multipar2025code]. Accordingly, we refer to this quantity as the pseudo-gain, following the discussion above. Naturally, the gain and pseudo-gain coincide whenever the vector prior mean $\boldsymbol{\mu}$ vanishes.
@@ Appendix B Optimality conditions for the MSL
for all $\boldsymbol{\tilde{\theta}}$. The minimum loss is then given by $\mathcal{L}_{\mathrm{min}}=\operatorname{Tr}(\Upsilon)$. A discrete version of this optimality condition exists for hypothesis testing [Yuen1970, Yuen1975], and a generalisation to convex functions has recently been proposed [Coutts2018].
@@
$\Upsilon=\Lambda+\tfrac{1}{2}L^{ij}\bigl(M_{\mathrm{opt},ji}\rho_{0}-\rho_{0}M_{\mathrm{opt},ji}-M_{\mathrm{opt},j}\bar{\rho}_{i}-\bar{\rho}_{j}M_{\mathrm{opt},i}\bigr).$ (112)
@@ C.1 Brief overview of monotone metrics
In this subsection we follow Hayashi2017c, but see also Ref. [Scandi2025] for a recent survey on the topic of monotone metrics. We define an inner product as follows:
@@ C.2 Posterior mean operators
Following Ref. [rubio2020bayesian], we now proceed to consider the scalar
@@
where $\theta_{\boldsymbol{u}}=\boldsymbol{u}^{T}\boldsymbol{\theta}=u^{i}\theta_{i}$ is a linear combination of parameters. For quantum systems, the outcome probability for a given value of the parameters is obtained from Born’s rule $p(x|\boldsymbol{\theta})=\operatorname{Tr}\left[M(x)\rho(\boldsymbol{\theta})\right]$, which leads to [rubio2020bayesian, Eq. (6)]
@@
A crucial observation from Ref. [rubio2020bayesian] is that the second term resembles the classical Fisher information and so it is possible to apply the Braunstein–Caves inequality [Braunstein1994] and obtain a matrix SPM lower bound on $\Sigma$. Here, we take a different route and use a more general family of inequalities which stem from the monotonicity of the inner products defined in the previous sections. This allows us to derive a whole family of matrix lower bounds, similarly to the case of local QET.
@@
where we have also introduced PM operators for each original parameter, satisfying the equations $E_{\rho_{0}}(L_{i})=\bar{\rho}_{i}$. Such PM operators are generalisation of the symmetric version introduced in the main text in terms of the map $E_{\rho,\mathrm{S}}$. For example, they also include the operators introduced in Ref. [Zhang2024g, Theorem 8], where they are called $\lambda$ logarithmic derivatives”.
@@
By choosing the map $E_{\rho,\mathrm{S}}$, this matrix becomes the real-valued SPM matrix introduced in Ref. [rubio2020bayesian]. However, for other choices, such as $E_{\rho,\mathrm{R}}$, this matrix is still Hermitian but complex-valued. Regardless of the particular choice of monotone metric, since the scalar inequality holds for an arbitrary choice of $\boldsymbol{u}$, we obtain the family of matrix lower bounds on the MSL matrix
@@
This is again a generalisation of the SPM matrix bound obtained in Ref. [rubio2020bayesian].
@@ C.3 Scalar bounds from matrix inequalities
We start from the matrix inequality in Eq. (139) and denote the right-hand side by the symbol $\mathcal{B}\coloneq\int\mathrm{d}\boldsymbol{\theta}\,p(\boldsymbol{\theta})\,\boldsymbol{\theta}\boldsymbol{\theta}^{\top}-\mathcal{K}$, where $\mathcal{K}\geq 0$ is a complex-valued Hermitian positive definite matrix. We now want to show that also the imaginary part of $\mathcal{K}$ sets constraints on the minimum MSL $\mathcal{L}$. This fact is well-known [holevo2011probabilistic]; for completeness, we provide a proof here, roughly following Ref. [demkowicz2020multi].
@@
where the absolute values come from choosing the $\pm$ sign corresponding to a positive quantity. The matrix $|\sqrt{L}\operatorname{Im}{\mathcal{B}}\sqrt{L}|$ is a Hermitian matrix with the same eigenbasis as $i\sqrt{L}\operatorname{Im}{\mathcal{B}}\sqrt{L}$, and with eigenvalues $|\lambda_{i}|$, while $\|A\|_{1}=\operatorname{tr}[\sqrt{A^{\dagger}A}]$ is the Schatten 1-norm (also known as trace norm or nuclear norm). [^8] Note that $\operatorname{tr}{|A|}=\|A\|_{1}$ only when $A$ is normal, as it is the case here. This inequality for the trace of the real part of a positive semidefinite complex matrix is known as the Belavkin–Grishanin inequality [Belavkin1973, Tsang2019].
@@
We finally note that the bound in Ref. [Zhang2024g, Theorem 8] in terms of the $\lambda$ logarithmic derivatives is also in the class of scalar bounds derived in this Appendix.

### v2

@@ Acknowledgements.
FA acknowledges fruitful discussions with Stefano Chessa, Rafał Demkowicz-Dobrzański, Vittorio Giovannetti and Ludovico Lami. FA acknowledges partial financial support from Marie Skłodowska-Curie Action EUHORIZON-MSCA-2021PF-01 (project QECANM, Grant No. 101068347). DB acknowledges support from the European Union’s Horizon Europe research and innovation program under the EPIQUE Project (Grant Agreement No. 101135288). JR acknowledges financial support from the Surrey Future Fellowship Programme. Code and data generated for this paper are openly available [167]. An open-source python package for evaluating the bounds discussed is also available on GitHub [111].
@@ Appendix A Prior loss and precision gain in Bayesian quantum estimation
where $\lambda$ was already defined in Eq. (15). Such an optimal prior loss is nothing but $\operatorname{tr}[L\Sigma_{p}]$, where $\Sigma_{p}$ is the covariance matrix of the multivariate prior probability $p(\boldsymbol{\theta})$ and we have introduced a lowercase trace of $d{\times}d$ matrices, distinct from the uppercase trace of operators acting on the physical Hilbert space of the system.
@@
Recent literature has referred to the quantity $L^{ij}\operatorname{Tr}[\rho_{0}S_{j}S_{i}]$ as an alternative definition of precision gain when the SPM bound is used as an optimiser for control parameters [35, 120]. From a computational standpoint, this definition is often more convenient, and it has likewise been adopted in our code [111]. Accordingly, we refer to this quantity as the pseudo-gain, following the discussion above. Naturally, the gain and pseudo-gain coincide whenever the vector prior mean $\boldsymbol{\mu}$ vanishes.
@@ Appendix B Optimality conditions for the MSL
for all $\boldsymbol{\tilde{\theta}}$. The minimum loss is then given by $\mathcal{L}_{\mathrm{min}}=\operatorname{Tr}(\Upsilon)$. A discrete version of this optimality condition exists for hypothesis testing [95, 96], and a generalisation to convex functions has recently been proposed [97].
@@
$\Upsilon=\Lambda+\tfrac{1}{2}L^{ij}\bigl(M_{\mathrm{opt},ji}\rho_{0}+\rho_{0}M_{\mathrm{opt},ji}-M_{\mathrm{opt},j}\bar{\rho}_{i}-\bar{\rho}_{j}M_{\mathrm{opt},i}\bigr).$ (111)
@@ C.1 Brief overview of monotone metrics
In this subsection we follow Hayashi [110, Ch. 6.1], but see also Ref. [135] for a recent survey on the topic of monotone metrics. We define an inner product as follows:
@@ C.2 Posterior mean operators
Following Ref. [40], we now proceed to consider the scalar
@@
where $\theta_{\boldsymbol{u}}=\boldsymbol{u}^{T}\boldsymbol{\theta}=u^{i}\theta_{i}$ is a linear combination of parameters. For quantum systems, the outcome probability for a given value of the parameters is obtained from Born’s rule $p(x|\boldsymbol{\theta})=\operatorname{Tr}\left[M(x)\rho(\boldsymbol{\theta})\right]$, which leads to [40, Eq. (6)]
@@
A crucial observation from Ref. [40] is that the second term resembles the classical Fisher information and so it is possible to apply the Braunstein–Caves inequality [69] and obtain a matrix SPM lower bound on $\Sigma$. Here, we take a different route and use a more general family of inequalities which stem from the monotonicity of the inner products defined in the previous sections. This allows us to derive a whole family of matrix lower bounds, similarly to the case of local QET.
@@
where we have also introduced PM operators for each original parameter, satisfying the equations $E_{\rho_{0}}(L_{i})=\bar{\rho}_{i}$. Such PM operators are generalisation of the symmetric version introduced in the main text in terms of the map $E_{\rho,\mathrm{S}}$. For example, they also include the operators introduced in Ref. [103, Theorem 8], where they are called $\lambda$ logarithmic derivatives”.
@@
By choosing the map $E_{\rho,\mathrm{S}}$, this matrix becomes the real-valued SPM matrix introduced in Ref. [40]. However, for other choices, such as $E_{\rho,\mathrm{R}}$, this matrix is still Hermitian but complex-valued. Regardless of the particular choice of monotone metric, since the scalar inequality holds for an arbitrary choice of $\boldsymbol{u}$, we obtain the family of matrix lower bounds on the MSL matrix
@@
This is again a generalisation of the SPM matrix bound obtained in Ref. [40].
@@ C.3 Scalar bounds from matrix inequalities
We start from the matrix inequality in Eq. (138) and denote the right-hand side by the symbol $\mathcal{B}\coloneq\int\mathrm{d}\boldsymbol{\theta}\,p(\boldsymbol{\theta})\,\boldsymbol{\theta}\boldsymbol{\theta}^{\top}-\mathcal{K}$, where $\mathcal{K}\geq 0$ is a complex-valued Hermitian positive definite matrix. We now want to show that also the imaginary part of $\mathcal{K}$ sets constraints on the minimum MSL $\mathcal{L}$. This fact is well-known [118]; for completeness, we provide a proof here, roughly following Ref. [64].
@@
where the absolute values come from choosing the $\pm$ sign corresponding to a positive quantity. The matrix $|\sqrt{L}\operatorname{Im}{\mathcal{B}}\sqrt{L}|$ is a Hermitian matrix with the same eigenbasis as $i\sqrt{L}\operatorname{Im}{\mathcal{B}}\sqrt{L}$, and with eigenvalues $|\lambda_{i}|$, while $\|A\|_{1}=\operatorname{tr}[\sqrt{A^{\dagger}A}]$ is the Schatten 1-norm (also known as trace norm or nuclear norm). [^8] Note that $\operatorname{tr}{|A|}=\|A\|_{1}$ only when $A$ is normal, as it is the case here. This inequality for the trace of the real part of a positive semidefinite complex matrix is known as the Belavkin–Grishanin inequality [100, 76].
@@
We finally note that the bound in Ref. [103, Theorem 8] in terms of the $\lambda$ logarithmic derivatives is also in the class of scalar bounds derived in this Appendix.
Appendix D Phase-dephasing multi-copy expressions
For the phase-dephasing model introduced in Sec. IV.2, the full single-copy zeroth and first state moments are
where $\operatorname{Ei}(z)$ denotes the exponential integral, $\Xi(z)=\displaystyle\sideset{{}_{3}}{{}_{3}}{\mathop{F}}((1,1,1),(2,2,2),z)$ is a generalised hypergeometric function, and $\gamma$ is Euler’s constant.
By virtue of the tensor product structure—although the moments themselves are not direct tensor products—the multi-copy state $\rho(\boldsymbol{\theta})=\varrho(\boldsymbol{\theta})^{\otimes n}$ consists of terms $\prod\limits_{j,k=1}^{d}[\varrho_{j,k}(\boldsymbol{\theta})]^{\kappa_{j,k}}$ for a $d$-dimensional system. For the multi-copy phase and dephasing example of Sec. IV.2, the pertinent integrals for the respective moments are given below. For $\rho_{0}$:
$\int\mathrm{d}\boldsymbol{\theta}\,\frac{[\varrho_{1,1}(\boldsymbol{\theta})]^{j}[\varrho_{2,2}(\boldsymbol{\theta})]^{n}}{\theta_{2}W_{1}\log W_{2}}$ $=\frac{1}{2^{j+n}},$ (146a)
$\int\mathrm{d}\boldsymbol{\theta}\,\frac{[\varrho_{1,1}(\boldsymbol{\theta})]^{j}[\varrho_{1,2}(\boldsymbol{\theta})]^{k}[\varrho_{2,1}(\boldsymbol{\theta})]^{k}[\varrho_{2,2}(\boldsymbol{\theta})]^{n}}{\theta_{2}W_{1}\log W_{2}}$ $=\frac{\left[\operatorname{Ei}(-2kW_{2})-\operatorname{Ei}\left(-\frac{2k}{W_{2}}\right)\right]}{2^{j+2k+n+1}\log W_{2}},$ (146b)
$\int\mathrm{d}\boldsymbol{\theta}\,\frac{[\varrho_{1,1}(\boldsymbol{\theta})]^{j}[\varrho_{1,2}(\boldsymbol{\theta})]^{k}[\varrho_{2,1}(\boldsymbol{\theta})]^{m}[\varrho_{2,2}(\boldsymbol{\theta})]^{n}}{\theta_{2}W_{1}\log W_{2}}$ $=\frac{\left[\operatorname{Ei}(-(k+m)W_{2})-\operatorname{Ei}\left(-\frac{(k+m)}{W_{2}}\right)\right]\sin\left(\frac{(k-m)W_{1}}{2}\right)}{2^{j+k+m+n}(k-m)W_{1}\log W_{2}},$ (146c)
where Eq. (146b) holds for $k>0$, and Eq. (146c) for $k\neq m$. For $\bar{\rho}_{1}$:
$\int\mathrm{d}\boldsymbol{\theta}\,\frac{[\varrho_{1,1}(\boldsymbol{\theta})]^{j}[\varrho_{1,2}(\boldsymbol{\theta})]^{k}[\varrho_{2,1}(\boldsymbol{\theta})]^{k}[\varrho_{2,2}(\boldsymbol{\theta})]^{n}\theta_{1}}{\theta_{2}W_{1}\log W_{2}}$ $=0,$ (147a)
$\int\mathrm{d}\boldsymbol{\theta}\,\frac{[\varrho_{1,1}(\boldsymbol{\theta})]^{j}[\varrho_{1,2}(\boldsymbol{\theta})]^{k}[\varrho_{2,1}(\boldsymbol{\theta})]^{m}[\varrho_{2,2}(\boldsymbol{\theta})]^{n}\theta_{1}}{\theta_{2}W_{1}\log W_{2}}$ $=\frac{\left[\operatorname{Ei}(-(k+m)W_{2})-\operatorname{Ei}\left(-\frac{(k+m)}{W_{2}}\right)\right]\sin\left(\frac{(k-m)W_{1}}{2}\right)}{2^{j+k+m+n}(k-m)W_{1}\log W_{2}},$ (147b)
where Eq. (147b) holds for $k\neq m$. For $\bar{\rho}_{2}$:
$\int\mathrm{d}\boldsymbol{\theta}\,\frac{[\varrho_{1,1}(\boldsymbol{\theta})]^{j}[\varrho_{2,2}(\boldsymbol{\theta})]^{n}\log\theta_{2}}{\theta_{2}W_{1}\log W_{2}}$ $=0,$ (148a)
$\int\mathrm{d}\boldsymbol{\theta}\,\frac{[\varrho_{1,1}(\boldsymbol{\theta})]^{j}[\varrho_{1,2}(\boldsymbol{\theta})]^{k}[\varrho_{2,1}(\boldsymbol{\theta})]^{k}[\varrho_{2,2}(\boldsymbol{\theta})]^{n}\log\theta_{2}}{\theta_{2}W_{1}\log W_{2}}$ $=\frac{1}{2^{j+2k+n+1}}\Bigg[\frac{k}{\log W_{2}}\left(W_{2}\Xi(-2kW_{2})-\frac{\Xi\left(-\frac{2k}{W_{2}}\right)}{W_{2}}\right)$ (148b)
$\mkern 128.0mu+\frac{\operatorname{Ei}\left(-2kW_{2}\right)+\operatorname{Ei}\left(-\frac{2k}{W_{2}}\right)}{2}-\log(2k)-\gamma\Bigg],$
$\int\mathrm{d}\boldsymbol{\theta}\,\frac{[\varrho_{1,1}(\boldsymbol{\theta})]^{j}[\varrho_{1,2}(\boldsymbol{\theta})]^{k}[\varrho_{2,1}(\boldsymbol{\theta})]^{m}[\varrho_{2,2}(\boldsymbol{\theta})]^{n}\log\theta_{2}}{\theta_{2}W_{1}\log W_{2}}$ $=\frac{\sin\left(\frac{(k-m)W_{1}}{2}\right)}{2^{j+k+m+n}W_{1}(k-m)}\Bigg[\frac{(k+m)}{2\log W_{2}}\left(W_{2}\Xi(-(k+m)W_{2})-\frac{\Xi\left(-\frac{k+m}{W_{2}}\right)}{W_{2}}\right)$ (148c)
$\mkern 128.0mu+\frac{\operatorname{Ei}\left(-(k+m)W_{2}\right)+\operatorname{Ei}\left(-\frac{k+m}{W_{2}}\right)}{2}-\log(k+m)-\gamma\Bigg],$
where Eq. (148b) holds for $k>0$, and Eq. (148c) for $k\neq m$.